# Arrow-Debreu Meets Kyle: Price Discovery for Derivatives

### v26

@@ Abstract
We analyze price discovery in a model where an agent has arbitrary private information regarding state probabilities and trades state-contingent claims. Our model integrates the foundational frameworks of Arrow and Debreu (1954) and Kyle (1985). In an equivalent options formulation, the informed agent has arbitrary information regarding an underlying asset’s payoff distribution and trades option portfolios. We characterize the informed demand, the price impact, and the information efficiency of prices. Our results explain practiced option trading strategies and the observed volatility smile across option strikes as equilibrium outcomes.
@@
Keywords: Contingent Claims; Price Discovery; Options; Straddle; Volatility Smile.
@@ 1 Introduction
A fundamental economic function of prices is to convey information. As informed agents trade to exploit their information regarding an asset’s expected payoff, their information is incorporated into the asset price. More broadly, agents often have information regarding an asset’s payoff distribution beyond its first moment—for example, its second moment (volatility)—which requires trading derivatives. Indeed, a principal utility of derivatives is to offer investors targeted exposure to various aspects of the underlying asset’s payoff distribution. Such derivative trades have long been standard practice, as seen in textbooks such as Hull (2015), which means that various information beyond the first moment is incorporated into derivatives prices.
@@
Despite their practical significance, existing literature lacks a comprehensive model that articulates how derivative markets systematically incorporate diverse information signals into prices, largely due to the challenges of modeling the complex information dynamics across these markets. This paper addresses this gap by developing a model that fully captures the nonlinearity of payoff distributions and the interdependence of trades and prices across derivatives. Integrating the foundational frameworks of Arrow and Debreu (1954) and Kyle (1985), our model offers robust characterizations of the informed demand, price impact, and information efficiency of prices across derivative markets.
@@
In our model, an informed trader with private information regarding the probabilities of future states trades state-contingent claims. In other words, the informed trader has information regarding the payoffs of Arrow-Debreu (AD) securities and trades in the AD markets. The model places no restrictions on the probability distribution across states or the type of information available to the informed trader.
@@
Replicating derivatives, such as options, using AD securities gives equivalent formulations of our model that account for the structural linkages across derivative securities. In an equivalent options formulation, the informed trader has information regarding an underlying asset’s payoff distribution and trades options, with no restriction on the payoff distribution or his information.
@@
The informed demand formula we obtain is simple and practical. It constructs the informed trader’s equilibrium portfolio of contingent claims using the possible distributions across states, assigning positive portfolio weight to the true distribution and negative weights to the others. When translated to options, the formula prescribes option strategies for trading on any aspect of the underlying payoff distribution. In particular, it rationalizes option trading strategies that are common in practice and widely discussed in empirical studies as equilibrium outcomes.
@@
When only the underlying asset is traded, our model reduces to the single-asset price discovery framework of Kyle (1985), where the equilibrium supply function is linear. The slope of the inverse supply function—known as “Kyle’s lambda”—determines the price impact and the extent to which demand reveals information about the asset’s payoff. This slope is proportional to the variation of the asset’s payoff. In the more general AD setting, demand for one security impacts prices across all markets based on the information it conveys about other securities. Extending the intuition behind Kyle’s lambda, we show that the equilibrium price impact between two securities is proportional to the covariance of their payoffs. Thus, Kyle’s lambda generalizes to an infinite-dimensional matrix of payoff covariances.
@@
In option markets, trades in one option generally have a price impact on the others, as they jointly convey information regarding the underlying payoff distribution. For example, trades for a put-call options pair, often used to trade on volatility, jointly convey information about the underlying payoff’s volatility. Our general characterization of cross-market price impact leads to new empirical measures and implications for this information spillover across derivatives, such as options across different strikes. Additionally, we show that the cross price impact of informed volatility trading can explain the “volatility smile”—an empirical regularity in option-implied volatilities that is central to reduced-form frameworks yet so far has lacked a microfounded explanation.
@@
The supply schedules for all the AD securities form a pricing kernel that prices all state-contingent claims under asymmetric information. In contrast to the one-dimensional linear supply function in the single-asset case, the pricing kernel is a high-dimensional nonlinear object that prices all derivatives simultaneously. While the pricing of derivatives has been extensively studied from the hedging perspective since Black and Scholes (1973), our contribution lies in characterizing their pricing under arbitrary private information, including providing a microfoundation for those derivatives trading patterns that the hedging motive alone cannot explain.
@@
Although empirical studies suggest that derivatives enhance market information efficiency (e.g., Roll et al. (2010)), theoretical explanations have remained limited. Our model provides a clear, intuitive explanation: The linear equilibrium of Kyle (1985) is not robust with respect to the introduction of derivative securities, as the absence of derivative markets constrains the informed trader to sub-optimal linear demands. Derivatives allow the informed trader to optimally reduce his demand for low-probability states (see Example 7.1). Since derivatives enable the informed trader to convey information that cannot be reflected in linear trades of the underlying asset, markets with derivatives achieve greater informational efficiency.
@@
In a complete market with derivatives, we show that the information efficiency of all prices is summarized in a single measure. Notably, this measure does not depend on the possible payoff distributions or noise trading levels across markets. Therefore, the extent of information incorporation into derivative prices is robust to specific informational details.
@@
Kyle (1985) provides a foundational model for single-asset price discovery. Extensions by Caballe and Krishnan (1994), Foster and Viswanathan (1996), and Back et al. (2000) analyze settings with multiple insiders possessing differential information across assets. Rochet and Vila (1994) relax the assumption of normal payoff distributions. Our work advances this strand of literature by analyzing the joint price discovery across derivative markets through interdependent trades in a general model.
@@
Empirical research has consistently shown that trades and prices across options markets reflect information beyond the expected payoff. For instance, Pan and Poteshman (2006) and Ni et al. (2008) find that option trading volumes predict future stock returns and volatility, while Goyal and Saretto (2009) show that mispricing in volatility can be exploited through straddles. Additional studies document that options order flows around specific events signal firm value and market expectations of future volatility—see Roll et al. (2009), Augustin et al. (2019) and Cao et al. (2005). Despite this empirical evidence for joint price formation across option markets, existing theoretical models of informed option trading have largely focused on comparing markets with or without a single option, as in Back (1993), Biais and Hillion (1994), Easley et al. (1998), and Collin-Dufresne et al. (2021). This leaves a theoretical gap in understanding how options trades and prices jointly incorporate broader information. We help bridge this gap by characterizating the mechanism of information incorporation across option markets.
@@
There have also been theoretical studies on non-strategic informed trading across multiple securities—see Admati (1985), Malamud (2015), and Chabakauri et al. (2022). The distinction of our strategic setting is that non-strategic models do not explain the practiced derivatives trading strategies or observed cross-strike empirical regularities. Additionally, we derive empirical implications showing how an informed trader’s private information (e.g., higher moments of an asset’s payoff) can be inferred from cross-security (e.g., cross-option) price impacts. This new insight is unique to the strategic setting.
@@
Many option pricing and financial econometric models have sought to capture the volatility smile by correcting mis-specifications within the Black and Scholes (1973) framework. Examples include the jump-diffusion models of Merton (1976)-type, [^1] See Bates (1991). the stochastic volatility models of Heston (1993)-type, [^2] See Bates (1996), Duffie et al. (2000), Britten-Jones and Neuberger (2000), Bates (2000), Aït-Sahalia and Kimmel (2007), Christoffersen et al. (2010), and Aït-Sahalia et al. (2021). and the local volatility models of Dupire (1994)-type. [^3] See Berestycki et al. (2002) and Carr and Cousot (2012). Our framework moves beyond reduced-form model mis-specifications to directly explain the strike structure of implied volatility through cross-strike information dynamics.
@@
The rest of the paper proceeds as follows. Section 2 demonstrates the basic intuition of our results in a simplified setting. Section 3 presents the general model. Section 4 shows that the market maker’s cross-market inferences are summarized by an intuitive sufficient statistic. Section 5 analyzes the informed trader’s cross-market price impact by decomposing his first-order condition. Section 6 establishes equilibrium through a canonical reduction of the trading game. While this canonical reduction is a key modeling innovation that makes our general model tractable, the broad intuition of the resulting equilibrium extends directly from Section 2 and can be gleaned through Example 6.8 without delving into the precise details. Section 7 discusses the characterization of price discovery for contingent claims—the informed demand, price impact, and information efficiency of prices. Section 8 explores the direct empirical implications of our results for price discovery across option markets. Section 9 concludes. Formal assumptions and proofs are provided in the Appendix.
@@
Figure 1a Conditional State Probabilities Solid dots - signal $s_{1}$ (high vol), hollow dots - signal $s_{2}$ (low vol).
Figure 1b
Insider Expected Profit Curve Equilibrium is characterized by the insider’s optimal portfolio size $\alpha^{*}$.
@@ 2 Basic Intuition
We begin with a simplified setting that encapsulates the basic economic intuition we bring out. In the general model, we will develop the robustness of this intuition—without restriction on the number of securities, payoff distributions, or private information.
There are two risk-neutral agents: the insider and the market maker. At $t=0$, the agents know that there are three possible $t=1$ states of the world, denoted as $x_{i}$, $i=1,2,3$, and they trade the corresponding AD securities.
At $t=0$, the insider privately observes one of two possible signals, $s_{1}$ or $s_{2}$. Conditional on signal $s_{1}$, the probability distribution over $t=1$ states is $(\frac{1}{2},0,\frac{1}{2})$. Conditional on $s_{2}$, the distribution is $(0,1,0)$. See Figure 1a. The market maker has uniform prior $\pi_{0}$ on signals.
@@
An intuitive trading strategy for the insider, conditional on observing $s_{1}$, is to buy securities $x_{1}$ and $x_{3}$ (where payoffs are positive) and sell security $x_{2}$ (where payoff is zero) for an equal $\alpha$ shares each. [^5] There are no leverage or short-selling constraints. Then, conditional on the insider observing $s_{1}$, the order flow received by the market maker across markets is $\omega=(\omega_{1},\omega_{2},\omega_{3})$ where
$\omega_{1}=\alpha+\varepsilon_{1},\;\omega_{2}=-\alpha+\varepsilon_{2},\;\omega_{3}=\alpha+\varepsilon_{3}.$
Similarly, suppose conditional on $s_{2}$ the insider buys security $x_{2}$ and sells securities $x_{1}$ and $x_{3}$ for $\alpha$ shares each.
Market Maker Inference
Given the insider’s trading strategy, the market maker’s posterior likelihood ratio conditional on order flow $\omega$ is
@@
The market maker sets prices based on the excess demand $\Delta(\omega)$ for $x_{1}$ and $x_{3}$ securities relative to $x_{2}$. If $\Delta(\omega)$ is high, he infers that the insider has likely observed $s_{1}$ and, in turn, raises $P_{1}$ and $P_{3}$ and lowers $P_{2}$, and vice versa. In other words, there is cross price impact, and it is determined by the covariances of security payoffs. Since the payoff covariance of securities $x_{1}$ and $x_{3}$ is $\frac{1}{8}$, their cross price impact is positive. The pairwise payoff covariance for the other two security pairs is $-\frac{1}{8}$, resulting in negative cross price impact. The covariance structure of payoffs determines the extent to which order flow for one security reveals information about others, therefore the cross price impact. This intuition extends directly to the general model developed in subsequent sections.
@@
Figure 1b illustrates the insider’s expected profit as a function of $\alpha$. If $\alpha\!=\!0$, the insider’s zero demand gives him zero profit, and the AD prices are $(\frac{1}{4},\frac{1}{2},\frac{1}{4})$—order flow contains no information, leaving the market maker’s posterior unchanged from his prior. If $\alpha\!\rightarrow\!\infty$, the market maker’s posterior probability of $s_{1}$ approaches one, causing the AD prices to converge to the true distribution $(\frac{1}{2},0,\frac{1}{2})$ and driving the insider’s profit to zero due to full information revelation.
@@
The insider’s optimal portfolio size $\alpha^{*}$ balances the insider’s marginal benefit from exploiting the true distribution $(\frac{1}{2},0,\frac{1}{2})$ against the marginal costs. These costs include the cost incurred at current prices, assuming no price impact, and the cost stemming from the cross-market price impact due to information revelation. His optimal portfolio size $\alpha^{*}$ characterizes equilibrium—this simple result will also hold in the general model. Allowing correlated noise trades, varying noise intensity, and non-uniform priors does not materially affect the results. [^6] Noise Trades Correlation Allowing noise trades to be correlated, while keeping their intensity constant, does not affect price impact. For example, suppose the correlation of noise trades for a pair of assets has the opposite sign as their payoff correlation. In that case, the insider would scale up his trading strategy to capture additional losses incurred by the noise traders trading against the payoffs. The market maker would shift his supply schedules accordingly, but the price impact at the margin would remain the same—this is similar to biased noise trades in the single-asset setting.  [^7] Noise Trading Intensity Suppose the noise trading intensity $(\sigma_{i})$ varies across markets, with $\epsilon_{i}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{i}^{2})$, $i=1,2,3$. Then $\Delta(\omega)$ becomes $\Delta(\omega)=2(\frac{\omega_{1}}{\sigma_{1}^{2}}+\frac{\omega_{3}}{\sigma_{3}^{2}}-\frac{\omega_{2}}{\sigma_{2}^{2}}).$ Thus, as expected, higher noise trading intensity reduces price impact, both within each market and across markets.  [^8] General Prior When the market maker has a general prior $(\pi_{0}(s_{1}),\pi_{0}(s_{2}))$, the posterior likelihood ratio (1) becomes $\frac{\pi_{1}(s_{1}|\omega)\pi_{0}(s_{1})}{\pi_{1}(s_{2}|\omega)\pi_{0}(s_{2})}=\frac{e^{\alpha\omega_{1}-\frac{1}{2}\alpha^{2}-\alpha\omega_{2}-\frac{1}{2}\alpha^{2}+\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}\pi_{0}(s_{1})}{e^{-\alpha\omega_{1}-\frac{1}{2}\alpha^{2}+\alpha\omega_{2}-\frac{1}{2}\alpha^{2}-\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}\pi_{0}(s_{2})}.$ The comparative statics for the prior follows intuitively. For example, a higher $\pi_{0}(s_{1})$ leads to greater cross price impact between securities $x_{1}$ and $x_{3}$—because it leads the market maker to place more emphasis on the co-movement of the orders for $x_{1}$ and $x_{3}$ in his inference.
@@
Suppose fewer than all three AD securities are available for trade. In that case, both the insider’s profit and the information efficiency of prices will decrease, with information efficiency measured by the posterior probability weight the market maker assigns to the true signal. Derivative markets allow the informed agents’s trades to better reflect their information, thus incorporating their private information into derivative prices. This is the basic economic logic motivating this paper.
@@
Suppose the states $x_{1}<x_{2}<x_{3}$ are the possible $t=1$ prices of an underlying asset whose $t=0$ price is $x_{2}$. In this case, $s_{1}$ is the “high volatility” signal, and $s_{2}$ the “low volatility” signal. The signal $s_{1}$ reveals that the asset price will move but without a specified direction while $s_{2}$ reveals that there will be no price movement. Security $x_{1}$ (resp. $x_{3}$) acts like a toy put (resp. call) option that pays off when the asset price declines (resp. rises). If the insider observes the high volatility signal $s_{1}$, he buys a toy put-call pair—a straddle—to capitalize on price movement regardless of direction. The actual straddle is a standard option strategy for trading on high volatility (see Hull (2015)) and is widely discussed in the empirical literature. [^9] See Coval and Shumway (2001), Pan and Poteshman (2006), Ni et al. (2008), Driessen et al. (2009), and Goyal and Saretto (2009). It and other common option strategies will be explained as equilibrium strategies in the general model.
The cross price impact between the put-call pair in the toy straddle is higher under the high volatility signal, since the insider’s demand for the put-call pair amplifies their cross-market information spillover. This means that information regarding the future volatility of the underlying payoff can be backed out from the cross price impact between the put-call pair in the straddle. In our general model, we systematically extend this insight to actual options, deriving empirical predictions about general aspects of the underlying asset’s payoff.

### v27

# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

@@ Abstract
We analyze price discovery in a model where an agent has arbitrary private information about state probabilities and trades state-contingent claims. Our model integrates the seminal frameworks of Arrow and Debreu (1954) and Kyle (1985). In an equivalent options formulation, the informed agent has arbitrary information about an underlying asset’s payoff distribution and trades option portfolios. We characterize the informed demand and information dynamics across markets. Our results offer the first equilibrium-based explanations for longstanding empirical practices and phenomena in option markets, including common trading patterns and the volatility smile across option strikes.
@@
Keywords: Contingent Claims, Price Discovery, Options, Straddle, Volatility Smile.
@@ 1 Introduction
A fundamental economic function of prices is to convey information. As informed agents trade to exploit their information about an asset’s expected payoff, that information is incorporated into the asset price. More broadly, agents often have information about an asset’s payoff beyond its mean, such as its volatility, which requires trading derivatives. Indeed, a principal utility of derivatives is to offer investors targeted exposure to various aspects of the underlying asset’s payoff distribution. Such derivative trading has long been standard practice, as covered in popular textbooks such as Hull (2015). This means that various information beyond the expected payoff is incorporated into derivatives prices.
@@
Yet, despite their practical significance, the existing literature lacks a comprehensive model that articulates how derivative markets systematically incorporate diverse information signals into prices. This gap arises largely due to the challenges of modeling the complex information dynamics across derivatives. We address this gap by developing a model that fully captures the nonlinearity of payoff distributions and the interdependence of trades and prices across derivatives. Integrating the foundational frameworks of Arrow and Debreu (1954) and Kyle (1985), we obtain general characterizations of the informed demand, price impact, and information efficiency of prices across derivative markets.
@@
In our model, an informed trader has private information about the probabilities of future states and trades state-contingent claims. In other words, the informed trader has information about the payoffs of Arrow-Debreu (AD) securities and trades in the AD markets. There is no restriction on the probability distribution across states or the type of information available to the informed trader.
@@
Replicating derivatives, such as options, using AD securities yields equivalent formulations of our model that account for the structural linkages across derivatives. In an equivalent options formulation, the informed trader has information about an underlying asset’s payoff distribution and trades options, with no restriction on the payoff distribution or his information.
@@
The informed demand formula we obtain is simple and practical. It constructs the informed trader’s contingent-claim portfolio by assigning positive portfolio weight to the true payoff distribution and negative weights to the others. Translated to options, the formula prescribes option strategies for trading on any aspect of the underlying payoff distribution. In particular, it rationalizes empirically observed option trading strategies as equilibrium outcomes—the first paper to do so.
@@
When only the underlying asset is traded, our model reduces to the single-asset price discovery framework of Kyle (1985), where the equilibrium supply function is linear. The slope of the inverse supply function—known as “Kyle’s lambda”—determines the price impact and the extent to which demand reveals information about the asset’s payoff. This slope is proportional to the variation of the asset’s payoff. In the more general AD setting, demand for one security impacts prices across all markets based on the information it conveys about other securities. Building on the intuition behind Kyle’s lambda, we show that the equilibrium price impact between two securities is proportional to the covariance of their payoffs. This generalizes Kyle’s lambda to an infinite-dimensional matrix of payoff covariances.
@@
In option markets, trades in one option typically affect the prices of others. This occurs because trades in options jointly convey information about the underlying payoff distribution. For example, a common volatility trade comprises a put-call options pair. Thus, trades in such a pair jointly convey information about volatility. Our general characterization of such cross-market spillovers introduces new empirical measures and predictions that link option cross-strike dynamics to different aspects of the underlying payoff. Moreover, we offer an intuitive equilibrium explanation for the “volatility smile,” an empirical regularity in option-implied volatilities that has been widely studied but limited to reduced-form models.
@@
The joint supply schedules for all AD securities form a pricing kernel for state-contingent claims under asymmetric information. In contrast to the one-dimensional linear supply function in the single-asset case, the pricing kernel is a high-dimensional nonlinear object that prices all derivatives simultaneously. While the pricing of derivatives has been extensively studied from the hedging perspective since Black and Scholes (1973), we extend beyond the hedging perspective to asymmetric information, providing a microfoundation for those trading and price patterns unexplained by hedging motives alone.
@@
Empirical evidence (e.g., Roll et al. (2010)) suggests derivatives improve market efficiency, but theoretical explanations are limited. Our model offers a clear explanation: the linear equilibrium of Kyle (1985) is not robust with respect to the introduction of derivative securities, as the absence of derivative markets constrains the informed trader to suboptimal linear demands. Derivatives allow the informed trader to optimally reduce his demand for low-probability states (see Example 7.1). By incorporating information that cannot be reflected in linear trades of the underlying asset, derivatives improve market efficiency.
@@
In a complete market with derivatives, we show that the information efficiency of all prices is summarized in a single measure. Notably, this measure does not depend on the possible payoff distributions or noise trading levels across markets. Thus, the informativeness of derivatives prices is robust to specific information details.
@@
Kyle (1985) is the seminal model for single-asset price discovery. Extensions by Caballe and Krishnan (1994), Foster and Viswanathan (1996), and Back et al. (2000) consider settings with multiple insiders holding differential information across assets. Rochet and Vila (1994) relax the assumption of normal payoff distributions. We expand the scope of this literature to encompass joint price discovery across contingent claim markets.
@@
Empirical research has consistently shown that trades and prices across options markets reflect information beyond the expected payoff. Pan and Poteshman (2006) and Ni et al. (2008) find that option trading volumes predict future stock returns and volatility, while Goyal and Saretto (2009) show that volatility mispricing can be exploited through option straddles. Other studies show that option order flows can signal firm value and market expectations of future volatility—see Roll et al. (2009), Augustin et al. (2019), and Cao et al. (2005). Despite this empirical evidence for joint price formation across option markets, theoretical models of informed option trading have been limited to no-option vs. single-option comparisons, as in Back (1993), Biais and Hillion (1994), Easley et al. (1998), and Collin-Dufresne et al. (2021). This creates a theoretical gap in understanding how trades and prices across options jointly reflect broader information. Our model helps bridge this gap; it provides a unified theoretical framework that consolidates empirical findings and suggests new avenues of empirical research.
@@
There have also been theoretical studies on non-strategic informed trading across multiple securities, such as Admati (1985), Malamud (2015), and Chabakauri et al. (2022). Non-strategic models do not explain the common derivatives trading strategies or observed cross-strike empirical regularities; moreover, Admati (1985) and Chabakauri et al. (2022) impose parametric assumptions on the payoff distribution while we do not. Additionally, we derive empirical predictions regarding how different payoff information (e.g., higher moments ) can be inferred from price impact across options. This new insight is unique to the strategic setting.
@@
Many option pricing and financial econometric models have sought to capture the volatility smile by correcting misspecifications in the Black and Scholes (1973) framework. Examples include the jump-diffusion models of Merton (1976)-type, [^1] See Bates (1991). the stochastic volatility models of Heston (1993)-type, [^2] See Bates (1996), Duffie et al. (2000), Britten-Jones and Neuberger (2000), Bates (2000), Aït-Sahalia and Kimmel (2007), Christoffersen et al. (2010), and Aït-Sahalia et al. (2021). and the local volatility models of Dupire (1994)-type. [^3] See Berestycki et al. (2002) and Carr and Cousot (2012). Our framework moves beyond reduced-form model misspecifications and directly explains the strike structure of implied volatility via cross-strike information dynamics.
@@
The rest of the paper proceeds as follows. Section 2 demonstrates the basic intuition of our results in a simplified setting. Section 3 presents the general model. Section 4 shows that the uninformed trader’s cross-market inferences are summarized by an intuitive sufficient statistic. Section 5 analyzes the informed trader’s cross-market price impact by decomposing his first-order condition. Section 6 establishes equilibrium through a canonical reduction of the trading game. While this canonical reduction is a key modeling innovation that makes our general model tractable, the broad intuition of the resulting equilibrium extends directly from Section 2 and can be understood from Example 6.7 without delving into the precise details. Section 7 discusses the characterization of price discovery for contingent claims—the informed demand, price impact, and information efficiency of prices. Section 8 explores the immediate empirical implications of our results for price discovery across option markets. Section 9 concludes. Formal assumptions and proofs are provided in the Appendix.
@@
(a) Conditional State Probabilities. Solid dots represent signal $s_{1}$ (high volatility), and hollow dots represent signal $s_{2}$ (low volatility).
Insider Expected Profit Curve. The insider’s optimal portfolio size $\alpha^{*}$ characterizes equilibrium.
@@ 2 Basic Intuition
We begin with a simplified setting that encapsulates the basic economic intuition we bring out. In the general model, we will develop the robustness of this intuition—without restriction on the number of securities, payoff distributions, or types of private information.
There are two risk-neutral agents: the insider and the market maker. At $t=0$, the agents know that there are three possible $t=1$ states, denoted $x_{i}$ for $i=1,2,3$, and trade the corresponding AD securities.
At $t=0$, the insider privately observes one of two possible signals, $s_{1}$ or $s_{2}$. Conditional on signal $s_{1}$, the probability distribution over $t=1$ states is $(\frac{1}{2},0,\frac{1}{2})$. Conditional on $s_{2}$, the distribution is $(0,1,0)$. See Figure 1a. The market maker has the uniform prior $\pi_{0}$ on signals.
After observing his private signal, the insider submits his demand (market orders) for AD securities to maximize his expected utility at $t=1$. Noise traders trade for exogenous reasons such as liquidity needs, and their trades across the AD markets are $\varepsilon_{i}\!\!\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\!\!\mathcal{N}(0,1)$ for $i=1,2,3$. The market maker acts as a competitive liquidity provider.
@@
An intuitive trading strategy for the insider, conditional on observing $s_{1}$, is to buy securities $x_{1}$ and $x_{3}$ (where payoffs are positive) and sell security $x_{2}$ (where payoff is zero) for some equal $\alpha$ shares each. [^5] There are no leverage or short-selling constraints. Then, conditional on the insider observing $s_{1}$, the order flow across markets received by the market maker is
$\omega=(\omega_{1},\omega_{2},\omega_{3}),\;\mbox{ where }\;\omega_{1}=\alpha+\varepsilon_{1},\;\omega_{2}=-\alpha+\varepsilon_{2},\;\omega_{3}=\alpha+\varepsilon_{3}.$
Similarly, conditional on $s_{2}$, the insider buys security $x_{2}$ and sells securities $x_{1}$ and $x_{3}$ for $\alpha$ shares each.
Market Maker’s Sufficient Statistic
Given the insider’s trading strategy, the market maker’s posterior likelihood ratio, conditional on order flow $\omega$, is
@@
Thus, the market maker’s sufficient statistic is the excess demand $\Delta(\omega)$ for $x_{1}$ and $x_{3}$ securities relative to $x_{2}$. If $\Delta(\omega)$ is high, he infers the insider likely observed $s_{1}$, leading him to raise $P_{1}$ and $P_{3}$ and lower $P_{2}$, and vice versa. In other words, there is cross price impact, and it is determined by the covariances of security payoffs. Since the payoff covariance of securities $x_{1}$ and $x_{3}$ is $\frac{1}{8}$, their cross price impact is positive. The pairwise payoff covariance for the other two security pairs is $-\frac{1}{8}$, resulting in negative cross price impact. The payoff covariances determine the extent to which order flow for one security reveals information about others, and thus determine the price impact across markets. This intuition extends directly to the general model developed in subsequent sections.
@@
Figure 1b illustrates the insider’s expected profit as a function of $\alpha$. If $\alpha\!=\!0$, the insider’s zero demand results in zero profit, and the AD prices are $(\frac{1}{4},\frac{1}{2},\frac{1}{4})$; since order flow contains no information, the market maker sets the prices based on his prior. If $\alpha\!\rightarrow\!\infty$, the market maker’s posterior probability of $s_{1}$ approaches one, which causes the AD prices to converge to the true distribution $(\frac{1}{2},0,\frac{1}{2})$, driving the insider’s profit to zero due to full information revelation.
@@
The insider’s optimal portfolio size $\alpha^{*}$ trades off the marginal payoff from exploiting the true distribution $(\frac{1}{2},0,\frac{1}{2})$ against the marginal costs. These costs include the cost incurred at current prices, assuming no price impact, and the cost stemming from the cross-market price impact due to information revelation. His optimal portfolio size $\alpha^{*}$ characterizes equilibrium—this simple result will extend to the general model. Allowing correlated noise trades, varying noise intensity, and non-uniform priors does not materially affect the results. [^6] Noise Trades Correlation Allowing noise trades to be correlated, while keeping their intensity constant, does not affect price impact. For example, suppose the correlation of noise trades for a pair of assets has an opposite sign to their payoff correlation. In that case, the insider would scale up his trading strategy to capture additional losses incurred by the noise traders trading against the payoffs. The market maker would shift his supply schedules accordingly, but the price impact at the margin would remain the same; this is similar to biased noise trades in the single-asset setting.  [^7] Noise Trading Intensity Suppose the noise trading intensity $(\sigma_{i})$ varies across markets, with $\epsilon_{i}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{i}^{2})$, $i=1,2,3$. Then $\Delta(\omega)$ becomes $\Delta(\omega)=2(\frac{\omega_{1}}{\sigma_{1}^{2}}+\frac{\omega_{3}}{\sigma_{3}^{2}}-\frac{\omega_{2}}{\sigma_{2}^{2}}).$ Thus, as expected, higher noise trading intensity reduces price impact within each market and across markets.  [^8] General Prior When the market maker has a general prior $(\pi_{0}(s_{1}),\pi_{0}(s_{2}))$, the posterior likelihood ratio (1) becomes $\frac{\pi_{1}(s_{1}|\omega)\pi_{0}(s_{1})}{\pi_{1}(s_{2}|\omega)\pi_{0}(s_{2})}=\frac{e^{\alpha\omega_{1}-\frac{1}{2}\alpha^{2}-\alpha\omega_{2}-\frac{1}{2}\alpha^{2}+\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}\pi_{0}(s_{1})}{e^{-\alpha\omega_{1}-\frac{1}{2}\alpha^{2}+\alpha\omega_{2}-\frac{1}{2}\alpha^{2}-\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}\pi_{0}(s_{2})}.$ The comparative statics for the prior follows intuitively. For example, a higher $\pi_{0}(s_{1})$ leads to greater cross price impact between securities $x_{1}$ and $x_{3}$—because it leads the market maker to place more emphasis on the co-movement of the orders for $x_{1}$ and $x_{3}$ in his inference.
@@
If fewer than all three AD securities are available for trade, both the insider’s profit and the information efficiency of prices decrease. Here, information efficiency is naturally measured by the posterior probability weight that the market maker assigns to the true signal. Derivative markets allow the informed agents’ trades to better reflect their information, thus incorporating their private information into prices. This is the basic economic logic motivating this paper.
@@
Suppose the states $x_{1}<x_{2}<x_{3}$ are the possible $t=1$ prices of an underlying asset whose $t=0$ price is $x_{2}$. In this case, $s_{1}$ is the “high volatility” signal, and $s_{2}$ the “low volatility” signal. The signal $s_{1}$ reveals that the asset price will move, but without a specified direction, whereas $s_{2}$ reveals that there will be no price movement. Security $x_{1}$ acts as a toy put option that pays off when the asset price declines, while $x_{3}$ acts as a call option that pays off when the asset price rises. If the insider observes the high volatility signal $s_{1}$, he buys a toy put-call pair—a straddle—to capitalize on price movement regardless of direction. The actual straddle is a standard option strategy for volatility trading (see Hull (2015)) and is widely discussed in the empirical literature. [^9] See Coval and Shumway (2001), Pan and Poteshman (2006), Ni et al. (2008), Driessen et al. (2009), and Goyal and Saretto (2009). This strategy, along with other common option strategies, will be explained as equilibrium strategies in the general model.
The cross price impact between the put-call pair in the toy straddle is greater under the high volatility signal, because the insider’s demand for the pair amplifies their cross-market information spillover. Therefore, information about the future volatility of the underlying payoff can be backed out from the cross price impact between the put-call pair in the straddle. In our general model, we systematically extend this insight to actual options and derive empirical predictions about general aspects of the underlying asset’s payoff.